1a. Arrange the compounds in order of decreasing PKa. Highest First.
a. CH3CH2OH
b. ClCH2CH2SH
c. CH3CH2SH

1b. Arrange the compounds in order of decreasing PKa. Highest First.
a. Cl2CHCH2OH
b. CH3CH2OH
c. ClCH2CH2OH
Answer #1
Concepts and reason

Strength of an acid is expressed in terms of Ka{{\rm{K}}_{\rm{a}}}(acid dissociation constant) value or pKa{\rm{p}}{{\rm{K}}_{\rm{a}}}(inverse logarithm value ofKa{{\rm{K}}_{\rm{a}}}) value.

The following factors determine the strength of an acidic compound.

1. The strength of an acid increases with increase in size of the atom to which donor hydrogen atom is attached in acidic compound.

2. Acidic strength increases with increase in electronegativity of the atom to which donor hydrogen atom attached.

3. Type of substituent and position of substituent from donor hydrogen atom will play role in determining acidic strength due to inductive effect.

Fundamentals

A proton (H+{{\rm{H}}^{\rm{ + }}}) donor is an acid. A compound which can donate a proton or hydrogen atom very easily is strongly acidic compound.

The structure of an acidic compound will affect its pKa{\rm{p}}{{\rm{K}}_{\rm{a}}}value.

Proton donating capacity of an acid is represented with pKa{\rm{p}}{{\rm{K}}_{\rm{a}}}value. A strong acid will have a small pKa{\rm{p}}{{\rm{K}}_{\rm{a}}}value and a weak acid will have a large pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} value.

Part 1.a

Compare the acid strength of CH3CH2OH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}} andCH3CH2SH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}. The compound CH3CH2OH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}is less acidic thanCH3CH2SH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}. The pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} value of CH3CH2OH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}} is higher than the pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} value ofCH3CH2SH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}, that ispKa(CH3CH2OH)>pKa(CH3CH2SH){\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}} \right) > {\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}} \right).

The compound CH3CH2SH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}is less acid thanClCH2CH2SH{\rm{ClC}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}. The pKa{\rm{p}}{{\rm{K}}_{\rm{a}}}value of CH3CH2SH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}is higher than the pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} value ofClCH2CH2SH{\rm{ClC}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}, that is pKa(CH3CH2SH)>pKa(ClCH2CH2SH){\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}} \right) > {\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{ClC}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}} \right).

Part 1.b

Compare the acid strength of CH3CH2OH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}} andClCH2CH2OH{\rm{ClC}}{{\rm{H}}_2}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}. The compound CH3CH2OH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}} is less acidic thanClCH2CH2OH{\rm{ClC}}{{\rm{H}}_2}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}. The pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} value of CH3CH2OH{\rm{C}}{{\rm{H}}_3}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}} is higher than the pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} value ofClCH2CH2OH{\rm{ClC}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}, that is pKa(CH3CH2OH)>pKa(ClCH2CH2OH){\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{C}}{{\rm{H}}_3}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}} \right) > {\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{ClC}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}} \right) .

The compoundClCH2CH2OH{\rm{ClC}}{{\rm{H}}_2}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}is less acid thanCl2CHCH2OH{\rm{C}}{{\rm{l}}_2}{\rm{CHC}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}. The pKa{\rm{p}}{{\rm{K}}_{\rm{a}}}value of ClCH2CH2OH{\rm{ClC}}{{\rm{H}}_2}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}is higher than the pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} value ofCl2CHCH2OH{\rm{C}}{{\rm{l}}_2}{\rm{CHC}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}, that is pKa(ClCH2CH2OH)>pKa(Cl2CHCH2OH){\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{ClC}}{{\rm{H}}_2}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}} \right) > {\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{C}}{{\rm{l}}_2}{\rm{CHC}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}} \right) .

[Part 1.b]

Ans: Part 1.a

Part 1.a

The decreasing order of pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} values of the given compounds ispKa(CH3CH2OH)>pKa(CH3CH2SH)>pKa(ClCH2CH2SH){\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}} \right) > {\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}} \right) > {\rm{p}}{{\rm{K}}_{\rm{a}}}\left( {{\rm{ClC}}{{\rm{H}}_{\rm{2}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{SH}}} \right).